OpenStudy (anonymous):

#6 Solve 2x - 4 = -7 justify each step #7 Your chorus holds a car wash they have $25.00 for making change at the end of the car wash they have $453.50 how much money did they make? #8 The rate to rent a certain truck is $55 per day and $0.20 per mile your family pays $80 to rent this truck for 1 day how many miles did your family drive? #9 Movie tickets for an adult and 3 children cost $20 and adult ticket cost $2 more than a child's ticket find the cost of the adult ticket?

OpenStudy (mathmale):

One question at a time, please, and before asking for help, would you please share what you have already thought or done towards solving a given problem? Which problem do you want to do in this post?

OpenStudy (anonymous):

#6

OpenStudy (anonymous):

2x-4=-7 4+2x=-7+4 addition 2x=-3 simplify 2x/2=-3/2 division x=-1.5 simplify